# Provable Benefit of Annealed Langevin Monte Carlo for Non-log-concave Sampling

### v1

@@ Abstract
We address the outstanding problem of sampling from an unnormalized density that may be non-log-concave and multimodal. To enhance the performance of simple Markov chain Monte Carlo (MCMC) methods, techniques of annealing type have been widely used. However, quantitative theoretical guarantees of these techniques are under-explored. This study takes a first step toward providing a non-asymptotic analysis of annealed MCMC. Specifically, we establish, for the first time, an oracle complexity of $\widetilde{O}\left(\frac{d\beta^{2}{\cal A}^{2}}{\varepsilon^{6}}\right)$ for simple annealed Langevin Monte Carlo algorithm to achieve $\varepsilon^{2}$ accuracy in Kullback-Leibler divergence to the target distribution $\pi\propto{\rm e}^{-V}$ on $\mathbb{R}^{d}$ with $\beta$-smooth potential $V$. Here, ${\cal A}$ represents the action of a curve of probability measures interpolating the target distribution $\pi$ and a readily sampleable distribution.
@@ 1 Introduction
We study the task of efficient sampling from a probability distribution $\pi\propto{\rm e}^{-V}$ on $\mathbb{R}^{d}$. This fundamental problem is pivotal across various fields including computational statistics [liu2004monte, robert2004monte, brooks2011handbook, owen2013monte], Bayesian inference [gelman2013bayesian], statistical physics [newman1999monte, landau2021a], and finance [dagpunar2007simulation], and has been extensively studied in the literature [chewi2022log]. The most common approach to this problem is Markov Chain Monte Carlo (MCMC), among which Langevin Monte Carlo (LMC) [durmus2019analysis, vempala2019rapid, chewi2022analysis, mousavi-hosseini2023towards] is a particularly popular choice. LMC can be understood as a time-discretization of a diffusion process, known as Langevin diffusion (LD), whose stationary distribution is the target distribution $\pi$, and has been attractive partly due to its robust performance despite conceptual simplicity.
@@
Although LMC and its variants converge rapidly when the target distribution $\pi$ is strongly log-concave or satisfies isoperimetric inequalities such as the log-Sobolev inequality (LSI) [durmus2019analysis, vempala2019rapid, chewi2022analysis], its effectiveness diminishes when dealing with target distributions that are not strongly log-concave or are multimodal, such as mixtures of Gaussians. In such scenarios, the sampler often becomes confined to a single mode, severely limiting its ability to explore the entire distribution effectively. This results in significant challenges in transitioning between modes, which can dramatically increase the mixing time, making it exponential in dimension $d$. Such limitations highlight the need for enhanced MCMC methodologies that can efficiently navigate the complex landscapes of multimodal distributions, thereby improving convergence rates and overall sampling efficiency.
@@
To address the challenges posed by multimodality, techniques around the notion of annealing have been widely employed [gelfand1990sampling, neal2001annealed]. The general philosophy involves constructing a sequence of intermediate distributions $\pi_{0},\pi_{1},...,\pi_{M}$ that bridge the gap between an easily samplable distribution $\pi_{0}$ (e.g., Gaussian or Dirac-like), and the target distribution $\pi_{M}=\pi$. The process starts with sampling from $\pi_{0}$ and progressively samples from each subsequent distribution until $\pi_{M}$ is reached. When $\pi_{i}$ and $\pi_{i+1}$ are close enough, approximate samples from $\pi_{i}$ can serve as a warm start for sampling from $\pi_{i+1}$, thereby facilitating this transition. Employing LMC within this framework gives rise to what is known as the annealed LMC algorithm, which is the focus of our study. Despite its empirical success [song2019generative, song2020improved, zilberstein2022annealed, zilberstein2024solving], a thorough theoretical understanding of annealed LMC, particularly its non-asymptotic complexity bounds, remains elusive.
@@
In this work, we take a first step toward providing a non-asymptotic analysis of annealed MCMC. By leveraging the Girsanov theorem to quantify the differences between the sampling dynamics and a reference process, we establish an upper bound of the error of annealed MCMC by the summation of two terms. One term is equal to the ratio of an action integral in the Wasserstein geometry induced by optimal transport and the duration of the process and diminishes when the annealing is sufficiently slow. Another term quantifies the discretization error in implementation. Our approach challenges the traditional point of view that annealed MCMC is a cascade of warm start. Analysis based on this viewpoint still requires log-concavity or isoperimetry. Instead, our theoretical analysis of the annealed Langevin-based MCMC advances the field by eliminating the reliance on assumptions related to log-concavity or isoperimetry. Our strategy represents a paradigm shift in analyzing annealed MCMC. Even though a fully polynomial complexity bound has not been achieved yet. We hope our strategy can inspire future work towards this ambitious goal.
@@
  1. 1.
    We discover a novel strategy to analyze the non-asymptotic complexity bound of annealed MCMC algorithms that can circumvent the need for log-concavity or isoperimetry.
  2. 2.
    In Section 4, we investigate the annealed LD, which involves running LD with a dynamically changing target distribution and derive a surprising bound on the time required to simulate the SDE for achieving $\varepsilon^{2}$-accuracy in KL divergence.
  3. 3.
    Building on the insights from the analysis of the continuous dynamic and accounting for discretization errors, we establish a non-asymptotic oracle complexity bound for annealed LMC in Section 5, which is applicable to a broad range of annealing schemes.
The quantitative results are summarized and compared to other sampling algorithms in Table 1. Notably, our approach operates under the least stringent assumptions and exhibits the most favorable $\varepsilon$-dependence among all isoperimetry-free sampling methods.
Table 1: Comparison of oracle complexities in terms of $d$, $\varepsilon$, and LSI constant for sampling from $\pi\propto{\rm e}^{-V}$. The notation “${\rm poly(\cdot)}$” denotes a polynomial function of the indicated parameters.
@@
[vempala2019rapid]
$C$-LSI Potential smooth $\varepsilon^{2}$, $\operatorname{KL}\left(\cdot\middle\|\pi\right)$ $\widetilde{O}\left(C^{2}d\varepsilon^{-2}\right)$
@@
[fan2023improved]
$C$-LSI Potential smooth $\varepsilon$, TV $\widetilde{O}\left(Cd^{1/2}\log\varepsilon^{-1}\right)$
@@
$\varepsilon$, TV $O\left(\operatorname{poly}\left(d,\varepsilon^{-1}\right)\right)$
@@
$\varepsilon$, TV $O\left(\operatorname{poly}\left(d\right){\rm e}^{\operatorname{poly}\left(\varepsilon^{-1}\right)}\right)$
@@
$\varepsilon$, TV+W2 ${\rm e}^{\widetilde{O}\left(d\right)O\left(\log\varepsilon^{-1}\right)}$
@@
/ Potential smooth $\varepsilon^{2}$, $\operatorname{KL}\left(\pi\middle\|\cdot\right)$ $\widetilde{O}\left(d{\cal A}(d)^{2}\varepsilon^{-6}\right)$
Related works and comparison. We provide a brief overview of the literature, mainly focusing on the algorithms for non-log-concave sampling and their theoretical analysis.
  • Samplers based on tempering. The fundamental concept of tempering involves sampling the system at various temperatures simultaneously: at higher temparatures, the distribution flattens, allowing particles to easily transition between modes, while at lower temperatures, particles can more effectively explore local structures. In simulated tempering [marinari1992simulated, woodard2009conditions], the system’s temperature is randomly switched, while in parallel tempering (also known as replica exchange) [swendsen1986replica, lee2023improved], the temperatures of two particles are swapped according to a specific rule. However, quantitative theoretical results for tempering are limited, and the existing results (e.g., [ge2018beyond, ge2018simulated, dong2022spectral]) apply primarily to certain special classes of non-log-concave distributions.
  • Samplers based on general diffusions. Inspired by score-based diffusion models [ho2020denoising, song2021denoising, song2021scorebased], recent advances have introduced sampling methods that reverse the Ornstein-Uhlenbeck process, as detailed in [huang2024reverse, huang2024faster, he2024zeroth]. These samplers exhibit reduced sensitivity to isoperimetric conditions, but rely on estimating score functions (gradients of log-density) via importance sampling, which poses significant challenges in high-dimensional settings. Concurrently, studies such as [zhang2022path, vargas2023denoising, vargas2024transport, richter2024improved] have employed neural networks to approximate unknown drift terms, enabling an SDE to transport a readily sampleable distribution to the target distribution. This approach has shown excellent performance in handling complex distributions, albeit at the expense of significant computational resources required for neural network training. In contrast, annealed LMC runs on a known interpolation of probability distributions, thus simplifying sampling by obviating the need for intensive score estimation or neural network training.
  • Non-asymptotic analysis for non-log-concave sampling. Drawing upon the stationary-point analysis in non-convex optimization, the seminal work [balasubramanian2022towards] characterizes the convergence of non-log-concave sampling via Fisher divergence. Subsequently, [cheng2023fast] applies this methodology to examine the local mixing of LMC. However, Fisher divergence is a relatively weak criterion compared to more commonly employed metrics such as total-variational distance or Wasserstein distances. In contrast, our study provides a convergence guarantee in terms of KL divergence, which implies convergence in total-variation distance and offers a stronger result.
@@ 2.1 Stochastic Differential Equations and Girsanov Theorem
A stochastic differential equation (SDE) $X=(X_{t})_{t\in[0,T]}$ is a stochastic process on $\Omega=C([0,T];\mathbb{R}^{d})$, the space of continuous functions from $[0,T]$ to $\mathbb{R}^{d}$. The dynamics of $X$ are typically represented by the equation ${\rm d}X_{t}=b(X_{t},t){\rm d}t+\sigma(X_{t},t){\rm d}B_{t}$, $t\in[0,T]$, where $(B_{t})_{t\in[0,T]}$ is a standard Brownian motion in $\mathbb{R}^{d}$. The path measure of $X$, denoted $\mathbb{P}^{X}$, characterizes the distribution of $X$ over $\Omega$ and is defined by $\mathbb{P}^{X}(A)=\operatorname{\mathbb{P}}\left(X\in A\right)$ for all measurable subset $A$ of $\Omega$. The following lemma, as a corollary of the Girsanov theorem [karatzas1991brownian, ustunel2013transformation, liptser2013statistics], provides a methodology for computing the KL divergence between two path measures and serves as a crucial technical tool in our proof.
@@
${\rm d}X_{t}=a_{t}(X_{t}){\rm d}t+\sqrt{2}{\rm d}B_{t},~{}X_{0}\sim\mu;\qquad{\rm d}Y_{t}=b_{t}(Y_{t}){\rm d}t+\sqrt{2}{\rm d}B_{t},~{}Y_{0}\sim\nu.$
@@
$\operatorname{KL}\left(\mathbb{P}^{X}\middle\|\mathbb{P}^{Y}\right)=\operatorname{KL}\left(\mu\middle\|\nu\right)+\frac{1}{4}\operatorname{\mathbb{E}}_{X\sim\mathbb{P}^{X}}\left[\int_{0}^{T}\left\|a_{t}(X_{t})-b_{t}(X_{t})\right\|^{2}{\rm d}t\right].$
@@ 2.2 Langevin Diffusion and Langevin Monte Carlo
${\rm d}X_{t}=-\nabla V(X_{t}){\rm d}t+\sqrt{2}{\rm d}B_{t},~{}t\in[0,\infty);~{}X_{0}\sim\mu_{0}.$ (1)
It is well-known that under mild conditions, $\pi$ is the unique stationary distribution this SDE, and when $\pi$ has good regularity properties, the marginal distribution of $X_{t}$ converges to $\pi$ as $t\to{+\infty}$, so we can sample from $\pi$ by simulating Equation 1 for a long time. However, in most of the cases, LD is intractable to simulate exactly, and the Euler-Maruyama discretization of Equation 1 leads to the Langevin Monte Carlo (LMC) algorithm. LMC with step size $h>0$ and target distribution $\pi\propto{\rm e}^{-V}$ is a Markov chain $\{X_{kh}\}_{k=0,1,...}$ constructed by iterating the following update rule:

### v2

@@ Abstract
We consider the outstanding problem of sampling from an unnormalized density that may be non-log-concave and multimodal. To enhance the performance of simple Markov chain Monte Carlo (MCMC) methods, techniques of annealing type have been widely used. However, quantitative theoretical guarantees of these techniques are under-explored. This study takes a first step toward providing a non-asymptotic analysis of annealed MCMC. Specifically, we establish, for the first time, an oracle complexity of $\widetilde{O}\left(\frac{d\beta^{2}{\cal A}^{2}}{\varepsilon^{6}}\right)$ for the simple annealed Langevin Monte Carlo algorithm to achieve $\varepsilon^{2}$ accuracy in Kullback-Leibler divergence to the target distribution $\pi\propto{\rm e}^{-V}$ on $\mathbb{R}^{d}$ with $\beta$-smooth potential $V$. Here, ${\cal A}$ represents the action of a curve of probability measures interpolating the target distribution $\pi$ and a readily sampleable distribution.
@@ 1 Introduction
We study the task of efficient sampling from a probability distribution $\pi\propto{\rm e}^{-V}$ on $\mathbb{R}^{d}$. This fundamental problem is pivotal across various fields including computational statistics (Liu, 2004; Brooks et al., 2011), Bayesian inference (Gelman et al., 2013), statistical physics (Newman & Barkema, 1999), and finance (Dagpunar, 2007), and has been extensively studied in the literature (Chewi, 2024). The most common approach to this problem is Markov Chain Monte Carlo (MCMC), among which Langevin Monte Carlo (LMC) (Durmus et al., 2019; Vempala & Wibisono, 2019; Chewi et al., 2022; Mousavi-Hosseini et al., 2023) is a particularly popular choice. LMC can be understood as a time-discretization of a diffusion process, known as Langevin diffusion (LD), whose stationary distribution is the target distribution $\pi$, and has been attractive partly due to its robust performance despite conceptual simplicity.
@@
Although LMC and its variants converge rapidly when the target distribution $\pi$ is strongly log-concave or satisfies isoperimetric inequalities such as the log-Sobolev inequality (LSI) (Durmus et al., 2019; Vempala & Wibisono, 2019; Chewi et al., 2022), its effectiveness diminishes when dealing with target distributions that are not strongly log-concave or are multimodal, such as mixtures of Gaussians. In such scenarios, the sampler often becomes confined to a single mode, severely limiting its ability to explore the entire distribution effectively. This results in significant challenges in transitioning between modes, which can dramatically increase the mixing time, making it exponential in problem parameters such as dimension, distance between modes, etc. (Dong & Tong, 2022; Ma et al., 2019). Such limitations highlight the need for enhanced MCMC methodologies that can efficiently navigate the complex landscapes of multimodal distributions, thereby improving convergence rates and overall sampling efficiency.
@@
To address the challenges posed by multimodality, techniques around the notion of annealing have been widely employed (Gelfand et al., 1990; Neal, 2001). The general philosophy involves constructing a sequence of intermediate distributions $\pi_{0},\pi_{1},...,\pi_{M}$ that bridge the gap between an easily samplable distribution $\pi_{0}$ (e.g., Gaussian or Dirac-like), and the target distribution $\pi_{M}=\pi$. The process starts with sampling from $\pi_{0}$ and progressively samples from each subsequent distribution until $\pi_{M}$ is reached. When $\pi_{i}$ and $\pi_{i+1}$ are close enough, approximate samples from $\pi_{i}$ can serve as a warm start for sampling from $\pi_{i+1}$, thereby facilitating this transition. Employing LMC within this framework gives rise to what is known as the annealed LMC algorithm, which is the focus of our study. Despite its empirical success (Song & Ermon, 2019; 2020; Zilberstein et al., 2023; 2024), a thorough theoretical understanding of annealed LMC, particularly its non-asymptotic complexity bounds, remains elusive.
@@
In this work, we take a first step toward developing a non-asymptotic analysis of annealed MCMC. Utilizing the Girsanov theorem to quantify the differences between the sampling dynamics and a reference process, we derive an upper bound on the error of annealed MCMC, which consists of two key terms. The first term is the ratio of an action functional in Wasserstein geometry, induced by optimal transport, to the duration of the process. This term decreases when the annealing schedule is sufficiently slow. The second term captures the discretization error inherent in practical implementations. Our approach challenges the traditional view that annealed MCMC is simply a series of warm starts. Analyses based on this perspective typically require assumptions of log-concavity or isoperimetric inequalities. In contrast, our theoretical framework for annealed MCMC dispenses with these assumptions, marking a significant shift in the understanding and analysis of annealed MCMC.
Contributions.
@@
  • We propose a novel strategy to analyze the non-asymptotic complexity bounds of annealed MCMC algorithms, bypassing the need for assumptions such as log-concavity or isoperimetry.
  • In Sec. 4, we investigate the annealed LD, which involves running LD with a dynamically changing target distribution. We derive a notable bound on the time required to simulate the SDE for achieving $\varepsilon^{2}$-accuracy in KL divergence.
  • Building on the insights from the analysis of the continuous dynamic and incorporating discretization errors, we establish a non-asymptotic oracle complexity bound for annealed LMC in Sec. 5, which is applicable to a wide range of annealing schemes.
Comparison of quantitative results.
The quantitative results are summarized and compared to other sampling algorithms in Tab. 1. For algorithms requiring isoperimetric assumptions, we include Langevin Monte Carlo (LMC, Vempala & Wibisono (2019)) and Proximal Sampler (PS, Fan et al. (2023)), which converges rapidly but do not have theoretical guarantees without isoperimetric assumptions. For isoperimetry-free samplers, we include Simulated Tempering Langevin Monte Carlo (STLMC, Ge et al. (2018b)), a tempering-based sampler converging rapidly for a specific family of non-log-concave distributions, and three algorithms inspired by score-based generative models: Reverse Diffusion Monte Carlo (RDMC, Huang et al. (2024a)), Recursive Score Diffusion-based Monte Carlo (RS-DMC, Huang et al. (2024b)), and Zeroth Order Diffusion-Monte Carlo (ZOD-MC, He et al. (2024)), which involve simulating the time-reversal of the Ornstein-Uhlenbeck (OU) process and require estimating the scores of the intermediate distributions. Notably, our approach operates under the least stringent assumptions and exhibits the most favorable $\varepsilon$-dependence among all isoperimetry-free sampling methods.
Table 1: Comparison of oracle complexities in terms of $d$, $\varepsilon$, and the LSI constant for sampling from $\pi\propto{\rm e}^{-V}$. “$\operatorname{poly}(\cdot)$” indicates a polynomial dependence on the specified parameters.
@@
$C$-LSI Potential smooth $\varepsilon^{2}$, $\operatorname{KL}(\cdot\|\pi)$ $\widetilde{O}(C^{2}d\varepsilon^{-2})$
@@
$C$-LSI Potential smooth $\varepsilon$, $\rm TV$ $\widetilde{O}(Cd^{1/2}\log\varepsilon^{-1})$
@@
$\varepsilon$, $\rm TV$ $O(\operatorname{poly}(d,\varepsilon^{-1}))$
@@
$\varepsilon$, $\rm TV$ $O(\operatorname{poly}(d){\rm e}^{\operatorname{poly}(\varepsilon^{-1})})$
RS-DMC
/ Potential smooth $\varepsilon^{2}$, $\operatorname{KL}(\pi\|\cdot)$ $\exp(O(\log^{3}d\varepsilon^{-2}))$
@@
$\varepsilon$, ${\rm TV}+{\rm W}_{2}$ $\exp(\widetilde{O}(d)O(\log\varepsilon^{-1}))$
@@
/ Potential smooth $\varepsilon^{2}$, $\operatorname{KL}(\pi\|\cdot)$ $\widetilde{O}(d{\cal A}(d)^{2}\varepsilon^{-6})$
Related works.
We provide a brief overview of the literature, mainly focusing on the algorithms for non-log-concave sampling and their theoretical analysis.
  1. 1.
    Samplers based on tempering. The fundamental concept of tempering involves sampling the system at various temperatures simultaneously: at higher temparatures, the distribution flattens, allowing particles to easily transition between modes, while at lower temperatures, particles can more effectively explore local structures. In simulated tempering (Marinari & Parisi, 1992; Woodard et al., 2009), the system’s temperature is randomly switched, while in parallel tempering (also known as replica exchange) (Swendsen & Wang, 1986; Lee & Shen, 2023), the temperatures of two particles are swapped according to a specific rule. However, quantitative theoretical results for tempering are limited, and the existing results (e.g., Ge et al. (2018a; b); Dong & Tong (2022)) apply primarily to certain special classes of non-log-concave distributions.
  2. 2.
    Samplers based on general diffusions. Inspired by score-based diffusion models (Ho et al., 2020; Song et al., 2021a; b), recent advances have introduced sampling methods that reverse the OU process, as detailed in Huang et al. (2024a; b); He et al. (2024). These samplers exhibit reduced sensitivity to isoperimetric conditions, but rely on estimating score functions (gradients of log-density) via importance sampling, which poses significant challenges in high-dimensional settings. Concurrently, studies such as Zhang & Chen (2022); Vargas et al. (2023; 2024); Richter & Berner (2024) have employed neural networks to approximate unknown drift terms, enabling an SDE to transport a readily sampleable distribution to the target distribution. This approach has shown excellent performance in handling complex distributions, albeit at the expense of significant computational resources required for neural network training. In contrast, annealed LMC runs on a known interpolation of probability distributions, thus simplifying sampling by obviating the need for intensive score estimation or neural network training.
  3. 3.
    Non-asymptotic analysis for non-log-concave sampling. Drawing upon the stationary-point analysis in non-convex optimization, the seminal work Balasubramanian et al. (2022) characterizes the convergence of non-log-concave sampling via Fisher divergence. Subsequently, Cheng et al. (2023) applies this methodology to examine the local mixing of LMC. However, Fisher divergence is a relatively weak criterion compared to more commonly employed metrics such as total-variational distance or Wasserstein distances. In contrast, our study provides a convergence guarantee in terms of KL divergence, which implies convergence in total-variation distance and offers a stronger result.
Notations and definitions.
@@ 2.1 Stochastic Differential Equations and Girsanov Theorem
A stochastic differential equation (SDE) $X=(X_{t})_{t\in[0,T]}$ is a stochastic process on $\Omega=C([0,T];\mathbb{R}^{d})$, the space of continuous functions from $[0,T]$ to $\mathbb{R}^{d}$. The dynamics of $X$ are typically represented by the equation $\mathrm{d}X_{t}=b_{t}(X)\mathrm{d}t+\sigma_{t}(X)\mathrm{d}B_{t}$, $t\in[0,T]$, where $(B_{t})_{t\in[0,T]}$ is a standard Brownian motion in $\mathbb{R}^{d}$, and $b_{t}(X)\in\mathbb{R}^{d}$, $\sigma_{t}(X)\in\mathbb{R}^{d\times d}$ depends on $(X_{s})_{s\in[0,t]}$. The path measure of $X$, denoted $\mathbb{P}^{X}$, characterizes the distribution of $X$ over $\Omega$ and is defined by $\mathbb{P}^{X}(A)=\operatorname{Pr}(X\in A)$ for all measurable subset $A$ of $\Omega$. The following lemma, as a corollary of the Girsanov theorem (Üstünel & Zakai, 2013), provides a methodology for computing the KL divergence between two path measures and serves as a crucial technical tool in our proof.
@@
$\mathrm{d}X_{t}=a_{t}(X)\mathrm{d}t+\sqrt{2}\mathrm{d}B_{t},~{}X_{0}\sim\mu;\qquad\mathrm{d}Y_{t}=b_{t}(Y)\mathrm{d}t+\sqrt{2}\mathrm{d}B_{t},~{}Y_{0}\sim\nu.$
@@
$\operatorname{KL}(\mathbb{P}^{X}\|\mathbb{P}^{Y})=\operatorname{KL}(\mu\|\nu)+\frac{1}{4}\mathbb{E}_{X\sim\mathbb{P}^{X}}\int_{0}^{T}\|a_{t}(X)-b_{t}(X)\|^{2}\mathrm{d}t.$
@@ 2.2 Langevin Diffusion and Langevin Monte Carlo
$\mathrm{d}X_{t}=-\nabla V(X_{t})\mathrm{d}t+\sqrt{2}\mathrm{d}B_{t},~{}t\in[0,\infty);~{}X_{0}\sim\mu_{0}.$ (1)
It is well-known that under mild conditions, $\pi$ is the unique stationary distribution this SDE, and when $\pi$ has good regularity properties, the marginal distribution of $X_{t}$ converges to $\pi$ as $t\to{+\infty}$, so we can sample from $\pi$ by simulating Eq. 1 for a long time. However, in most of the cases, LD is intractable to simulate exactly, and the Euler-Maruyama discretization of Eq. 1 leads to the Langevin Monte Carlo (LMC) algorithm. LMC with step size $h>0$ and target distribution $\pi\propto{\rm e}^{-V}$ is a Markov chain $\{X_{kh}\}_{k=0,1,...}$ constructed by iterating the following update rule:
@@
2.3 Isoperimetric Inequalities
A probability measure $\pi$ on $\mathbb{R}^{d}$ satisfies a log-Sobolev inequality (LSI) with constant $C$, or $C$-LSI, if for all $f\in C^{1}(\mathbb{R}^{d})$ with $\mathbb{E}_{\pi}{f^{2}}>0$,
$\mathbb{E}_{\pi}{f^{2}\log\frac{f^{2}}{\mathbb{E}_{\pi}{f^{2}}}}\leq 2C\mathbb{E}_{\pi}{\|\nabla f\|^{2}}.$
A probability measure $\pi$ on $\mathbb{R}^{d}$ satisfies a Poincaré inequality (PI) with constant $C$, or $C$-PI, if for all $f\in C^{1}(\mathbb{R}^{d})$,
$\operatorname{Var}_{\pi}f\leq C\mathbb{E}_{\pi}{\|\nabla f\|^{2}}.$
It is worth noting that $\alpha$-strongly-log-concave distributions satisfy $\frac{1}{\alpha}$-LSI, and $C$-LSI implies $C$-PI (Bakry et al., 2014). It is established in Vempala & Wibisono (2019) that when $\pi\propto{\rm e}^{-V}$ satisfies $C$-LSI, the LD converges exponentially fast in KL divergence; furthermore, when the potential $V$ is $\beta$-smooth, the LMC also converges exponentially with a bias that vanishes when the step size approaches 00.